Suppose the Payroll tax reduction for middle-income households has been extended in the amount of $200 billion for the remainder of 2018. Assuming the MPC for that income group of households is 0.8 and also assuming that other things stay the same, the increase in GDP under this proposed extension of tax break is expected to increase by (ΔY) _______          1 pt a. $2,000 billion

  1. $1,000 billion
  2. $800 billion
  3. $500 billion

Hint: Need to use tax multiplier and it is a negative tax.
Answer #1

Solution: $1,000 billion

Working:

Multiplier = 1 / (1-MPC) = 1/MPS

= 1 / (1-0.8) = 5

Therefore an increase in GDP under this proposed extension of tax break is expected to raise ΔY by 1,000 billion (=200 billion * 5)